OpenStudy (anonymous):

how would you solve 6^(x^2+14x-38)=36^(6x-7) how would you solve 6^(x^2+14x-38)=36^(6x-7) @Mathematics

OpenStudy (jamesj):

36 = 6^2 thus 36^(6x -7) = 6^(12x - 14)

OpenStudy (aravindg):

james help meee

OpenStudy (jamesj):

So now you know that 6^(x^2+14x-38)=6^(12x - 14) and that can only be the case if x^2 + 14x - 38 = 12x - 14.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (alfie):

$$6^{x^{2}+14x-38}=36^{6x-7}$$ $$6^{x^{2}+14x-38}=6^{2(6x-7)}$$ $$x^{2}+14x\ -38\ =\ 12x\ -14$$